OpenStudy (anonymous):

what is the sequence in this pattern 1/10, 1/5, 3/10, 2/5

OpenStudy (lasttccasey):

\[x+\frac{1}{10}\]is your pattern. x being your previous number.. \[\frac{1}{10},\frac{2}{10},\frac{3}{10},\frac{4}{10}\]just simplify the even ones of course

OpenStudy (anonymous):

well ty

OpenStudy (anonymous):

how did they get 1/10 and 3/10 12345678910

OpenStudy (anonymous):

i still dont get it

OpenStudy (lasttccasey):

Your welcome, to give you a better sequence... \[\frac{1}{10},\frac{1}{5},\frac{3}{10},\frac{2}{5},\frac{1}{2}\]is basically the same as\[\frac{1}{10},\frac{2}{10},\frac{3}{10},\frac{4}{10},\frac{5}{10},\] just simplified

OpenStudy (anonymous):

oh ty

OpenStudy (lasttccasey):

Your welcome, understand it?

OpenStudy (anonymous):

yes